# Enhancing LLM-based Search Agents via Contribution Weighted Group Relative Policy Optimization

### 15 Apr 2026 v1

@@ 5.3 Supervision Decomposition
wo-reasoning 23.5 55.5 31.0 34.0 28.5 12.25 31.5 28.56
Table 4: Ablation on contribution supervision in CW-GRPO. We remove supervision on retrieval utility (wo-retrieval) or reasoning correctness (wo-reasoning), allowing a search round to contribute if only one condition is satisfied. Dropping either signal leads to inferior performance compared to full CW-GRPO, demonstrating that both are necessary for effective credit assignment in search agents.
@@ 5.4 Training Dynamics
As shown in Figure 3, the optimization remains stable across training, with no signs of divergence or collapse. The Exact Match score steadily improves, while the number of task rounds stays well-controlled and converges to a stable range. In addition, both retrieval utility and reasoning correctness exhibit consistent upward trends, indicating progressive improvement in the agent’s search behavior. These results show that CW-GRPO achieves stable optimization while improving both final accuracy and the quality of the search process.
We further analyze the effect of the conjunctive gate mechanism during training. Among all search rounds in successful trajectories, approximately 13.8% are filtered out (i.e., assigned zero contribution), indicating that the gate meaningfully reshapes the distribution of learning signals rather than acting as a near-identity transformation.
@@ 5.5 Case Study
Figure 4 compares the search trajectories of Qwen3-8B before and after CW-GRPO training on a factual question regarding John I. Jenkins’s tenure start. The untrained model issues only a single query fails to distinguish between the election year and the actual commencement of his term, leading to an incorrect answer based on this temporal confusion.
In contrast, the CW-GRPO trained model adopts a qualitatively different strategy by issuing multiple parallel queries in the first round. This expanded search coverage allows the agent to recall a broader set of documents, including explicit evidence that Jenkins’s term began in 2005. This case illustrates that CW-GRPO not only improves retrieval utility through parallelization but also strengthens the agent’s resistance to distractor evidence. By learning to prioritize direct answer-bearing evidence over superficially relevant but misleading signals, the model achieves more robust and grounded search-based decision making.
@@ 6 Discussion: Generalization Across Tasks and Models
CW-GRPO is motivated by the general principle of redistributing outcome-derived advantages according to per-round contribution estimates. This principle is broadly applicable beyond the specific experimental settings considered in this work.
6.0.0.1 Task generalization.
CW-GRPO can be applied to multi-round agentic tasks where the contributions of intermediate rounds can be reasonably estimated. In long-horizon decision-making environments(e.g., Webshop [yao2022webshop]), individual actions should correspond to incremental progress toward latent subgoals. This enables a judge to assess state improvements at each round and facilitates per-round credit assignment.
The framework also extends to single-round but structurally decomposable reasoning tasks. In settings such as mathematical reasoning(e.g., MATH [hendrycks2021measuringmathematicalproblemsolving]), model outputs typically consist of multi-step derivations. When a trajectory fails, the error can often be localized to specific steps; these steps can then be assigned higher responsibility for failure, enabling a CW-GRPO variant based on step-level attribution.
However, CW-GRPO is less suitable for tasks where per-step contribution is ill-defined or inherently unstable. This includes highly subjective generation tasks (e.g., open-ended creative writing) and perceptual generation tasks (e.g., image synthesis), where holistic judgments dominate and assigning reliable credit to intermediate steps via low-variance signals is challenging.
6.0.0.2 Model generalization.
CW-GRPO is model-agnostic and does not rely on a specific backbone architecture. Nevertheless, stronger base models tend to be more sensitive to the accuracy of per-round supervision, placing higher demands on the quality of the judge or rubric design, which may limit the practical gains of the method.
@@ 7 Conclusion
Finally, CW-GRPO relies on an external LLM judge for contribution estimation. While we find the associated computational cost to be manageable in our setting (see Appendix C), this design still introduces additional inference cost and dependency on judge quality. Exploring lighter-weight supervision signals (e.g., distilled or heuristic judges) may further improve scalability, especially in more resource-constrained or latency-sensitive scenarios.
@@ Appendix D Prompts
Figure 5 illustrates the prompt used by the LLM judge, which specifies the rubric for assessing each search round in terms of retrieval utility and reasoning correctness, producing the binary contribution signals used in CW-GRPO.

### 20 Apr 2026 v2

@@ 5.3 Supervision Decomposition
These results highlight the necessity of joint supervision: effective search requires both acquiring novel information and correctly reasoning over it.
@@
wo-reasoning 23.5 55.5 31.0 34.0 28.5 12.3 31.5 28.56
Table 4: Ablation on contribution supervision in CW-GRPO. Removing either retrieval utility or reasoning correctness supervision degrades performance, showing both are necessary for effective credit assignment.
@@ 5.4 Training Dynamics
As shown in Figure 3, the optimization remains stable across training, with no signs of collapse. The Exact Match score steadily improves, accompanied by consistent upward trends in both retrieval utility and reasoning correctness, indicating progressive improvement in the agent’s search behavior. These results show that CW-GRPO achieves stable optimization while improving both final accuracy and the quality of the search process.
We further analyze the effect of the conjunctive gate mechanism during training. Among all search rounds in successful trajectories, approximately 13.84% (2.2k out of 15.8k) are assigned zero contribution, indicating that the gate meaningfully reshapes the distribution of learning signals rather than acting as a near-identity transformation.
@@ 5.5 Case Study
Figure 4 compares the search trajectories of Qwen3-8B before and after CW-GRPO training on a factual question regarding John I. Jenkins’s tenure start. The untrained model issues only a single query and fails to distinguish between the election year and the actual commencement of his term, leading to an incorrect answer based on this temporal confusion.
In contrast, the CW-GRPO-trained model adopts a qualitatively different strategy by issuing multiple parallel queries in the first round. This expanded search coverage allows the agent to recall a broader set of documents, including explicit evidence that Jenkins’s term began in 2005. This case illustrates that CW-GRPO not only improves retrieval utility through parallelization but also strengthens the agent’s resistance to distracting evidence. By learning to prioritize direct answer-bearing evidence over superficially relevant but misleading signals, the model achieves more robust and grounded search-based decision making.
@@ 6 Discussion: Generalization Across Tasks and Models
CW-GRPO is motivated by the general principle of redistributing outcome-derived advantages according to round-level contribution estimates. This principle is broadly applicable beyond the specific experimental settings considered in this work.
Task generalization. CW-GRPO can be applied to multi-round agentic tasks where the contributions of intermediate rounds can be reasonably estimated. In long-horizon decision-making environments(e.g., Webshop [Yao et al., 2022]), individual actions should correspond to incremental progress toward latent subgoals. This enables a judge to assess state improvements at each round and facilitates round-level credit assignment.
The framework also extends to single-round but structurally decomposable reasoning tasks. In settings such as mathematical reasoning(e.g., MATH [Hendrycks et al., 2021]), model outputs typically consist of multi-step derivations. When a trajectory fails, the error can often be localized to specific steps; these steps can then be assigned higher responsibility for failure, enabling a CW-GRPO variant based on step-level attribution.
However, CW-GRPO is less suitable for tasks where process contribution is ill-defined or inherently unstable. This includes highly subjective generation tasks (e.g., open-ended creative writing) and perceptual generation tasks (e.g., image synthesis), where holistic judgments dominate and assigning reliable credit to intermediate steps via low-variance signals is challenging.
Model generalization. CW-GRPO is model-agnostic and does not rely on a specific backbone architecture. Nevertheless, stronger base models tend to be more sensitive to the accuracy of round-level supervision, placing higher demands on the quality of the judge or rubric design, which may limit the practical gains of the method.
@@ 7 Conclusion
Finally, CW-GRPO relies on an external LLM judge for contribution estimation. While we find the associated computational cost to be manageable in our setting (see Appendix C), this design still introduces additional inference cost and dependency on judge quality. Exploring lighter-weight supervision signals (e.g., heuristic judges) may further improve scalability, especially in more resource-constrained or latency-sensitive scenarios.
@@ Appendix D Prompts
We present the detailed prompts that guide the behavior of our LLM judge and search agent in this section.
Figure 5 illustrates the prompt used by the LLM judge, which specifies the rubrics for assessing each search round in terms of retrieval utility and reasoning correctness, producing the binary contribution signals used in CW-GRPO.